# Label-Efficient LiDAR Panoptic Segmentation

### v1

@@ III-C Primal Label Generation
Clustering and Correction: For each scene, we extract point clusters separately for both partitions and using HDBSCAN [44] (step 3). Furthermore, all points that were not assigned to any cluster by HDBSCAN are collected in an additional cluster . We reassign all points in to their respective nearest cluster in using a -nearest neighbors classifier. Intuitively, the geometry-based clusters should approximate semantic objects, i.e., all points within a cluster should receive the same pseudo-label (step 4). We further define a set of rare semantic classes that are underrepresented in the primal labels, e.g., construction vehicle. For each cluster, we compute the frequency of each semantic class. If the most frequent class is void and the frequency is greater than a threshold , we assign all points in the cluster to void. If the frequency of a rare class is above a threshold , we use this class as the pseudo-label of all points within the cluster. Otherwise, we use the cluster’s most frequent label:
@@ IV-A Dataset and Evaluation Metrics
To evaluate our method, we require a dataset that provides synchronized LiDAR scans and surround-view images. Although this setup is common in mobile robotics, ground truth panoptic annotations are rarely available for both modalities in the same dataset. While LiDAR annotations are required for evaluation, a small number of image annotations are required to train the label-efficient image segmentation network. The most widely used benchmarks [21, 22, 4, 34, 36, 35] for LiDAR panoptic segmentation are Panoptic nuScenes [16] and SemanticKITTI [42]. Neither of these datasets provides annotated surround-view images. However, the nuImages dataset [41] was recorded in the same domain as nuScenes and contains labels for one stuff class and the same ten thing classes as available in Panoptic nuScenes. To match the convention of Panoptic nuScenes, we manually add the labels for five additional stuff classes to a small () subset of images. Following previous work [14, 15], we carefully select these images to capture all semantic categories. We then use these human-annotated images to train PASTEL [15], followed by employing L3PS to generate panoptic point cloud pseudo-labels for the nuScenes dataset. Throughout this section, we report results on the val split using the panoptic quality (PQ) and mean intersection over union (mIoU) metrics.
@@ V Conclusion
In this work, we addressed the challenge of reducing the annotation effort for LiDAR panoptic segmentation. We proposed the novel method L3PS, which transfers recent label reduction techniques from the image domain to the point cloud domain. Our L3PS approach consists of two steps, namely (1) leveraging a highly label-efficient method for visual panoptic segmentation to generate image pseudo-labels that are projected onto 3D LiDAR scans and (2) refining these pseudo-labels based on spatial 3D information. We show that the refined point pseudo-labels can be used to train an off-the-shelf network for LiDAR panoptic segmentation achieving competitive performance to various baselines while shifting the annotation effort to the less complex image space. We believe that our work represents an important step towards cost-efficient point cloud segmentation contributing to robotic 3D perception. Future research could explore integrating our approach with unsupervised domain adaptation techniques and further advancing point cloud representation learning, drawing parallels to recent progress in the computer vision domain.

### v2

@@ III-C Primal Label Generation
Clustering and Correction: For each scene, we extract point clusters separately for both partitions and using HDBSCAN [44] (step 3). Furthermore, all points that were not assigned to any cluster by HDBSCAN are collected in an additional cluster . We reassign all points in to their respective nearest cluster in using a -nearest neighbors classifier. Intuitively, the geometry-based clusters should approximate semantic objects, i.e., all points within a cluster should receive the same pseudo-label (step 4). We further define a set of rare semantic classes that are underrepresented in the primal labels, e.g., construction vehicle. For each cluster, we compute the frequency of each semantic class. If the most frequent class is void and the frequency is greater than a threshold , we assign all points in the cluster to void. If the frequency of one or multiple rare classes is above a threshold , we use the most frequent rare class as the pseudo-label of all points within the cluster. Otherwise, we use the cluster’s most frequent label:
@@ IV-A Dataset and Evaluation Metrics
To evaluate our method, we require a dataset that provides synchronized LiDAR scans and surround-view images. Although this setup is common in mobile robotics, ground truth panoptic annotations are rarely available for both modalities in the same dataset. While LiDAR annotations are required for evaluation, a small number of image annotations are required to train the label-efficient image segmentation network. The most widely used benchmarks [21, 22, 4, 34, 36, 35] for LiDAR panoptic segmentation are Panoptic nuScenes [16] and SemanticKITTI [42]. Neither of these datasets provides annotated surround-view images. However, the nuImages dataset [41] was recorded in the same domain as nuScenes and contains labels for one stuff class and the same ten thing classes as available in Panoptic nuScenes. To match the convention of Panoptic nuScenes, we manually add the labels for the five missing stuff classes to a small () subset of images.111These annotations are available with our code. Following previous work [14, 15], we carefully select these images to capture all semantic categories. We then use these human-annotated images to train PASTEL [15], followed by employing L3PS to generate panoptic point cloud pseudo-labels for the nuScenes dataset. Throughout this section, we report results on the val split using the panoptic quality (PQ) and mean intersection over union (mIoU) metrics.
@@ V Conclusion
In this work, we addressed the challenge of reducing the annotation effort for LiDAR panoptic segmentation. We proposed the novel method L3PS, which transfers recent label reduction techniques from the image domain to the point cloud domain. Our L3PS approach consists of two steps, namely (1) leveraging a highly label-efficient method for visual panoptic segmentation to generate image pseudo-labels that are projected onto 3D LiDAR scans and (2) refining these pseudo-labels based on spatial 3D information. We show that the refined point pseudo-labels can be used to train an off-the-shelf network for LiDAR panoptic segmentation, achieving competitive performance compared to various baselines while shifting the annotation effort to the less complex image space. We believe that our work represents an important step towards cost-efficient point cloud segmentation, contributing to robotic 3D perception. Future research could explore integrating our approach with unsupervised domain adaptation techniques and further advancing point cloud representation learning, drawing parallels to recent progress in the computer vision domain.